OpenStudy (anonymous):

Paisley has x number of dimes and y number of nickels for a total of 18 coins. Knowing that a dime is worth ten cents and a nickel is worth five cents, she wrote the system of equations below because she had a total of $1.40. Is her system correct? If not, then why? x + y = 18 0.1x + 0.05y = 1.40

OpenStudy (nikato):

What do you think?

zepdrix (zepdrix):

Look at the first equation, it might easier to start with that one. x= number of dimes y= number of nickels Let's write that in place of our equation:\[\Large\rm (number~of~dimes)+(number~of~nickels)=18(coins)\]Does this first equation match up which what our paragraph is telling us?

zepdrix (zepdrix):

Comeon Keriiiiii >.< Gimme some feedbaccccck. Don't just stare at the screen :3 lol

OpenStudy (anonymous):

haha, omg i'm sorry!

zepdrix (zepdrix):

First equation match the information? Or does it cause a problem?

OpenStudy (anonymous):

i think sso

zepdrix (zepdrix):

So the first sentence is telling us that there are 18 coins, consisting of both nickels and dimes. Ok good it looks like that equation matches the information. Second one is a little trickier. We know that a dime is worth 10 cents. But we want to write cents in terms of `dollars`. So 10 cents written in terms of dollars looks like .10, ok? A dime is a 10th of a dollar, or .10 of a dollar.

zepdrix (zepdrix):

So then .10 multiplying x is 10cents times the number of dimes. So this is telling us the `values of the dimes`

zepdrix (zepdrix):

Let's write that into our equation and see if it makes sense,\[\Large\rm 0.10x + 0.05y = 1.40\]\[\large\rm (value~of~dimes)+ (value~of~nickels) = 1.40(total~value~of~coins)\]

zepdrix (zepdrix):

Does this one match up with our information?

OpenStudy (aric200):

the world may never know

zepdrix (zepdrix):

hehe :P

OpenStudy (anonymous):

yeh

zepdrix (zepdrix):

Mmmk good! So it looks like her system is correct!

OpenStudy (anonymous):

Thank you. >///<

OpenStudy (aric200):

No problem

OpenStudy (aric200):

Because I helped so much

OpenStudy (anonymous):

haha xD